OpenStudy (anonymous):

Factor the polynomial. x^9+y^9

OpenStudy (dumbcow):

sum of cubes \[a^3 +b^3 = (a+b)(a^2 -ab+b^2)\]

Directrix (directrix):

x^9 + y^9 is the sum of two cubes and will factor. x^9 + y^9 = (x^3)^3 + (y^3)^3